Commander Report – 03 May 2018

Commander Report
Anima Patil-Sabale
Mission Sol 11 (Earth Day May 3rd, 2018)

With only a day left in our mission, we are a bit disappointed we missed the opportunity for our exploration EVA’s two days in a row. We hope to make the best of our EVA window tomorrow.

Weather on Mars has been unpredictable, keeping us inbound for an entire sol.

We had some interstellar visitors today, a group of young, future space explorers.
Our sol was spent interacting with them, giving them a tour of the Mars Desert Research Station facilities here and addressing their curiosities about space travel and living on Mars.

Instilling inspiration and arousing curiosity in young minds is a mission all astronauts love, so it was a great experience enjoyed by the crew and our visitors.

As we approach our last sol on Mars, we work on summarizing all the work that we have done during this mission and look forward to making the best of our last sol here on Mars tomorrow. For now, we spend a relaxed evening enjoying a Thai style dinner and maybe catching up on our last movie on Mars.

Astronomy Report – 03 May 2018

Astronomy Report

Name: Eric Shear

Crew: 193
Date: May 3, 2018

Sky Conditions: Clear until about noon, heavy clouds in the afternoon.

Wind Conditions: Low

Observation Start Time: 1100

Observation End Time: 1500


No pictures were taken with the camera today, because we hosted a group of schoolkids on a tour of the entire campus, including Musk. The astronomer set up the telescope to track the sun and each kid was allowed to look through it. There were two subgroups of kids, one at around 1130 and the other at around 1430.

After the kids left, the astronomer closed up the observatory because of threatening rain.

Objects Viewed: Sun

Problems Encountered:

Clouds began to obscure the sun during the second viewing.

Sol Summary Report – 03 May 2018

Sol 11

Summary Title: Sol Close to the End

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Nominal

Sol Activity Summary: The entire day was spent touring around a class of 30 middle schoolers from Colorado. They viewed the Hab, the Green Hab, the Science Dome, and the Musk Observatory. They children had several good questions that our crew was able to discuss with them. They were excited about the possibilities of travelling to mars and becoming astronauts in the future! The crew enjoyed the time with the class but also long for the exploration of EVAs as they were cancelled yesterday due to rain. We intend on completing our final crew EVAs of this mission tomorrow.

Look Ahead Plan:

Tomorrow is our final day in simulation on Mars. We are looking forward to two EVAs, getting the hab ready for handover on Saturday and finishing our final reports.

Anomalies in work: None to report today, but the mud has now returned.

Weather: Partly cloudy in AM, cloudy in the PM with rain.

Crew Physical Status: Nominal


No EVAs were completed today due to the touring class from Colorado.

Science Report – 03 May 2018

1. Emergency EVAC EVA: No EVAs occurred today due to visit from an elementary school class.

2. Sunspot and Solar Flare Monitoring: No pictures were taken with the camera today, because we hosted a group of school kids on a tour of the entire campus.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to visit from an elementary school class.

4. GPS Route Measurements: No EVAs occurred today due to visit from an elementary school class.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the last two days of the mission. The study is continuing to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to visit from an elementary school class.

Operations Report – 03 May 2018

SOL: 11

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger.
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 9:00 pm to 9:15 am.

Solar— SOC 19 % (Before generator is run at night)

Diesel – 65 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 30 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134569.7

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): 350.1

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.1

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Had visitors to HAB today, spent most time giving a tour

Summary of internet: We’re out, same as normal.
Summary of suits and radios: One piece suit #10 is repaired. Still work to do on two piece suits.

Summary of Hab operations: Muddily nominal

Summary of GreenHab operations: Nominal; watered twice

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Children ate lunch inside RAM, CO secured after their visit.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Nothing yet.

GreenHab Report – 03 May 2018

Date: 3May2018

Environmental control: Ambient with door open in the afternoon

Shade cloth: on

Fan: Briefly

Average temperatures: 39.0℃

Low: 15.2℃

High: 43.5℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 8 Gallons

Time(s) of watering for crops: Twice: at 0915 and again at 1615

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: The plants are very happy with their care, but are eager to try some of this ‘Starbucks’ they’ve been hearing so much about, so if you would please pass some this way…

Support/supplies needed: None

EVA Report – 03 May 2018

Sol 11 Crew 193 EVA Summary
Earth-Date 3MAY2018

EVAs _
There were no EVAs planned today, as we had visitors this morning that lasted into the afternoon. After they left we had some rain, which precluded any contingency EVA that we had planned.

We did have an Engineering EVA this morning, in which our CO used an ATV to make contact with the visitors from another planet.

Copyright © The Mars Society. All rights reserved. | Main Site